# PortraitTalk: Towards Customizable One-Shot Audio-to-Talking Face Generation

### v1

@@ Introduction
Our model integrates three modalities for talking face generation, offering output flexibility. Figures 5 and 1 illustrate various examples, including identity interpolation, expression editing, and modifications of style, age, gender, and background. This multimodal integration substantially enhances the customization capability of PortraitTalk, facilitating the generation of emotionally accurate talking faces. However, PortraitTalk occasionally exhibits degraded visual quality and lip synchronization, especially under intense emotions. Additionally, unintended elements such as hands might appear when modifying the style, attributable to the exclusive reliance on the CLIP text encoder, without a specialized training dataset for style changes. These issues highlight the need for enhancements to achieve an optimal balance between expressiveness and realism of the outputs. Exhibition of Artifacts is provided in supplementary material.
Conclusion
@@
We introduced PortraitTalk, a customizable one-shot audio-to-talking face generation framework, comprising IdentityNet and AnimateNet. PortraitTalk preserves the identity and enhances the temporal coherence in video generation. It not only delivers high-fidelity audio-lip synchronization but also offers flexible customization capability via text prompts. This enables the creation of expressive talking faces across varied styles and emotions without retraining. Both quantitative and qualitative assessments confirm the superiority of our model, affirming its effectiveness and practical applicability in real-world scenarios.

### v2

@@ IV-I Applications and Limitations
PortraitTalk offers extensive customization capabilities for talking face generation through the integration of visual, textual, and audio modalities. Fig. 8 and Fig. 1 illustrate a range of applications, including identity interpolation, expression editing, and modifications of style, age, gender, and background. This multimodal conditioning significantly enhances PortraitTalk’s ability to generate emotionally expressive and visually coherent talking faces. However, certain limitations persist. As shown in Fig. 9, PortraitTalk occasionally exhibits degraded visual quality and reduced lip synchronization accuracy, particularly when synthesizing intense emotional expressions. In addition, unintended elements, such as hands, may appear during style modifications, likely due to the reliance on CLIP-based text embeddings without dedicated training for complex style alterations. These observations highlight the need for future improvements to achieve a better balance between customization flexibility, expressiveness, and output realism.
Figure 9: Limitations of the PortraitTalk model. While PortraitTalk supports a wide range of controllable generation features, certain failure cases persist. These include the appearance of unintended elements, changes in color contrast, and alterations in hairstyle that do not align with the intended prompt or reference. Such artifacts are more likely to occur in scenarios involving complex style manipulations or strong emotional expressions.
V Conclusion
@@
We introduced PortraitTalk, a robust and customizable audio-to-talking face generation framework, comprising IdentityNet and AnimateNet. PortraitTalk preserves the identity and enhances the temporal coherence in video generation. It not only delivers high-fidelity audio-lip synchronization but also offers flexible customization capability via text prompts. This enables the creation of expressive talking faces across varied styles and emotions without retraining. Both quantitative and qualitative assessments confirm the superiority of our model, affirming its effectiveness and practical applicability in real-world scenarios.